The problem that "A" can Solve a given problem is 4/5 that "B" can solve is 2/3 and "C" can solve 3/4 . If All 3 try, compute the probability will be solved.
Answer by greenestamps(13200)



P(at least one solves the problem)

= 1 - P(none of them solves the problem)

= 1 - (1-(4/5))*(1-(2/3))*(1-(3/4))

= 1 - (1/5)(1/3)(1/4)

= 1 - 1/60

= 59/60